# Large Language Model-Brained GUI Agents: A Survey

### v9

@@ 9.6 Computer Agent Benchmarks
Evaluating GUI agents on desktop computers involves diverse applications and complex workflows. Several benchmarks have been developed to assess agents’ capabilities in these environments, each addressing specific challenges and advancing the field. We overview benchmarks for computer GUI agents in Table XXVIII.
@@
TABLE XXIX: Overview of cross-platform GUI agent benchmarks.
Benchmark Platform Year Highlight Data Size Metric Measurement Link
VisualAgent Bench [298] Web, Android, Game, Virtual Embodied. 2024 First benchmark designed for visual foundation agents across GUI and multimodal tasks, focusing on vision-centric interactions in Android, web, and game environments. 4,482 trajectories Task Success Rate Text Match https://github.com/THUDM/VisualAgentBench/
SPR Benchmark [365] Mobile, Web, and Operating Systems 2024 Evaluates GUI screen readers’ ability to describe both content and layout information Includes 650 screenshots annotated with 1,500 target points and regions Task Success Rate, Efficiency Score Text Match, Element Match /
AgentStudio [385] Windows, Linux, macOS 2024 Open toolkit for creating and benchmarking general-purpose virtual agents, supporting complex interactions across diverse software applications. NA Step Success Rate Action Match, State Information and Image Match https://computer-agents.github.io/agent-studio/
CRAB [369] Linux, Android 2024 Cross-environment benchmark evaluating agents across mobile and desktop devices, using a graph-based evaluation method to handle multiple correct paths and task flexibility. 120 tasks Step Success Rate, Effiency Score Action Match https://github.com/crab-benchmark
ScreenSpot [25] iOS, Android, macOS, Windows, Web. 2024 Vision-based GUI benchmark with pre-trained GUI grounding, assessing agents’ ability to interact with GUI elements across mobile, desktop, and web platforms using only screenshots. 1,200 instructions Step Success Rate Action Match https://github.com/njucckevin/SeeClick
@@ 9.7 Cross-Platform Agent Benchmarks
To develop GUI agents capable of operating across multiple platforms, cross-platform benchmarks are essential. These benchmarks challenge agents to adapt to different environments and interfaces, evaluating their versatility and robustness. We provide an overview of benchmarks for cross-platform GUI agents in Tables XXIX.
@@
Complementing this, CRAB [369] introduces an innovative benchmark by evaluating multimodal language model agents in cross-environment interactions. It uniquely supports seamless multi-device task execution, evaluating agents in scenarios where tasks span both Ubuntu Linux and Android environments. By introducing a graph-based evaluation method that breaks down tasks into sub-goals and accommodates multiple correct paths to completion, CRAB provides a nuanced assessment of planning, decision-making, and adaptability. Metrics such as Completion Ratio, Execution Efficiency, Cost Efficiency, and Success Rate offer comprehensive insights into agent performance.
@@ 9.8 Takeaways
    Towards More Interactive and Realistic Environments: Recent advancements in GUI agent benchmarking emphasize the transition from synthetic scenarios to more interactive and realistic environments. This shift is evident in the use of simulators, Docker containers, and real-world applications to create "live" environments that better mimic genuine user interactions. Such environments not only provide a more accurate assessment of agent capabilities but also pose new challenges in terms of performance and robustness.
@@
    Emphasis on Safety, Privacy, and Compliance: There is a notable trend towards evaluating GUI agents on safety, privacy, and compliance metrics. These considerations are increasingly important as agents are integrated into sensitive and regulated domains. Encouraging this trend will help ensure that agents not only perform tasks effectively but also adhere to necessary legal and ethical standards. Future benchmarks should continue to expand on these dimensions, incorporating evaluations that reflect real-world compliance and data security requirements.
The landscape of GUI agent benchmarking is rapidly evolving to meet the demands of increasingly complex and interactive environments. By embracing cross-platform evaluations, fostering human-like interactions, addressing scalability challenges, and prioritizing safety and compliance, the community can pave the way for the next generation of sophisticated GUI agents. Continued innovation and collaboration will be essential in refining benchmarks to ensure they accurately capture the multifaceted capabilities of modern agents, ultimately leading to more intuitive and effective human-computer interactions.
@@ 10.1 GUI Testing
TABLE XXX: Overview of GUI-testing with LLM-powered GUI agents (Part I).
@@
TABLE XXXI: Overview of GUI-testing with LLM-powered GUI agents (Part II).
@@
Figure 25 and illustrates the use of an LLM-powered GUI agent to test font size adjustment on Windows OS. With only a natural language test case description, the agent autonomously performs the testing by executing UI operations, navigating through the settings menu, and leveraging its screen understanding capabilities to verify the final outcome of font size adjustment. This approach dramatically reduces the effort required for human or script-based testing. Next, we detail the GUI testing works powered by GUI agents, and first provide an overview Tables XXX and XXXI.
@@ 10.2 Virtual Assistants
TABLE XXXII: Overview of virtual assistants with LLM-powered GUI agents (Part I).
@@
TABLE XXXIII: Overview of virtual assistants with LLM-powered GUI agents (Part II).
@@
To explore more real-world applications of virtual assistants powered by GUI agents, we provide an overview of advancements across research, open-source initiatives, and production-level applications, as summarized in Table XXXII and XXXIII.
@@ 11.1 Privacy Concerns
LLM-powered GUI agents often require access to sensitive user data, including screenshots, personal credentials, interaction logs, and confidential documents, which may need to be transmitted to remote servers for processing. This cloud-based deployment raises significant privacy risks, such as data breaches, unauthorized access, and potential misuse of personal information [452, 453, 454]. For instance, consider an agent that automates email management. To sort emails or compose responses, the agent needs access to the user’s email content, which may include sensitive information. Transmitting this data to a cloud server for processing could expose the user to privacy risks. Users may be hesitant to adopt such agents due to concerns over data security and privacy violations [455, 456].
Potential Solutions: To mitigate privacy concerns, future research should focus on enabling on-device inference, where the language model operates directly on the user’s device without uploading personal data [457, 458]. Achieving this requires advancements in model compression techniques [459], on-device optimization [460], and efficient inference algorithms [461] to accommodate the computational limitations of user devices. Additionally, implementing privacy-preserving techniques like federated learning [462], differential privacy [463], and homomorphic encryption [464] can enhance data security while allowing the model to learn from user data.
Furthermore, developers of GUI agents should collaborate with privacy policymakers to ensure that user data and privacy are appropriately protected [465]. They should make the data handling processes transparent to users, clearly informing them about what data are being transmitted and how they are used, and obtain explicit user consent [466].
@@ 11.2 Latency, Performance, and Resource Constraints
Since GUI agents rely on LLMs, which are large models with substantial computational demands, this can result in high latency and slow response times, negatively affecting the user experience [467]. In time-sensitive applications, delays in action execution can lead to user frustration or even system failure. This issue becomes more pronounced in long-term tasks, where latency can accumulate at every step, exacerbating the problem. Furthermore, when using on-device inference with resource-constrained devices, the situation can be even more severe [468]. For example, a mobile app incorporating an LLM-powered agent may experience slow performance or excessive battery drain, detracting from the user experience [469].
@@ 11.3 Safety and Reliability
The ability of GUI agents to perform real-world actions within software environments could introduce safety [476, 454] and reliability [477] concerns. Erroneous actions could result in unintended consequences, such as data corruption, application crashes, or security vulnerabilities [478]. The probabilistic nature of LLM outputs means that agents may occasionally generate incorrect, inconsistent or hallucinated actions [479, 480]. For example, an agent tasked with automating financial transactions might misinterpret a command and transfer funds to the wrong account, leading to financial losses. The agent may also be vulnerable to black-box attacks, which could compromise its functionality and security [481]. Additionally, integrating GUI agents into existing software ecosystems involves compatibility issues and security considerations, and may encounter resistance from users who distrust automated systems.
@@ 11.4 Human-Agent Interaction
When using a GUI agent, any interruption or interaction by the user in the environment may interfere with the agent. Resolving such conflicts and designing the relationship between the human user and the GUI agent becomes challenging [493, 494]. In addition, users may provide vague or ambiguous requests, leading agents to misunderstand the intended task. Furthermore, agents may encounter situations where they lack sufficient information or face unexpected obstacles, or need confirmation from the user for certain action [19]. Determining when and how the agent should seek human assistance or clarification is crucial for effective collaboration [495]. This is common in daily usage when the agent does not have sufficient context [496]. Furthermore, users may need to intervene in the agent’s behavior if it does not meet their expectations.
As illustrated in the fabricated example shown in Figure 27, when the agent is tasked with sending an email to Tom, it engages in several steps to ensure privacy, accuracy, and user intent. First, the agent requests the user to manually log in, preserving sensitive credentials like usernames and passwords. Next, when multiple matches are found for the recipient “Tom”, the agent resolves the ambiguity by prompting the user to select the correct contact. Finally, before sending the email, the agent seeks confirmation from the user, acknowledging that sending an email is a sensitive action since it cannot be retracted [19]. This seemingly simple request demonstrates the complexity of human-agent interaction, requiring careful design to handle privacy, ambiguity, and confirmation efficiently [497]. It highlights the challenges and intricacies involved in developing robust and user-friendly GUI agents.
@@
Potential Solutions: Emphasizing user-centered design [498] principles can address user needs and concerns, providing options for customization and control over the agent’s behavior [495]. Equipping agents with the ability to engage in clarification dialogues when user instructions are unclear can enhance task accuracy [499]. Natural language understanding components can detect ambiguity and prompt users for additional information. For instance, the agent could ask, “There are two contacts named John. Do you mean John Smith or John Doe?” Incorporating human-in-the-loop systems allows for human intervention during task execution, enabling users to guide or correct the agent’s decisions when necessary [500]. Developing adaptive interaction models that facilitate seamless collaboration between humans and agents is essential. Additionally, providing transparency and explainability in the agent’s reasoning processes can build user trust and improve cooperation [501, 502].

### v11

@@ 9.6 Computer Agent Benchmarks
Evaluating GUI agents on desktop computers involves diverse applications and complex workflows. Several benchmarks have been developed to assess agents’ capabilities in these environments, each addressing specific challenges and advancing the field. We overview benchmarks for computer GUI agents in Table XXXVIII.
@@
Computer Agent Arena [451] presents a new paradigm for benchmarking LLM-based GUI agents through live, user-configured desktop environments. Unlike traditional static datasets, it provides an interactive cloud-based infrastructure where agents are evaluated on tasks spanning web browsing, programming, and productivity using real applications like Google Docs, VSCode, and Slack. Its innovation lies in using head-to-head agent comparisons, human judgment, and Elo-based ranking to evaluate general-purpose digital agents in realistic settings. The benchmark supports Windows and Ubuntu, with MacOS support planned, and allows customized task scenarios with diverse software and website setups. By enabling crowdsourced evaluations and planning open-source releases, it fosters community-driven improvements and robust comparisons.
@@
TABLE XXXIX: Overview of cross-platform GUI agent benchmarks.
Benchmark Platform Year Live Highlight Data Size Metric Measurement Link
VisualAgent Bench [368] Web, Android, Game, Virtual Embodied 2024 Yes First benchmark designed for visual foundation agents across GUI and multimodal tasks, focusing on vision-centric interactions in Android, web, and game environments. 4,482 trajectories Task Success Rate Text Match https://github.com/THUDM/VisualAgentBench/
SPR Benchmark [455] Mobile, Web, Operating Systems 2024 Yes Evaluates GUI screen readers’ ability to describe both content and layout information 650 screenshots annotated with 1,500 target points and regions Task Success Rate, Efficiency Score Text Match, Element Match /
AgentStudio [238] Windows, Linux, macOS 2024 Yes Open toolkit for creating and benchmarking general-purpose virtual agents, supporting complex interactions across diverse software applications. NA Step Success Rate Action Match, State Information, Image Match https://computer-agents.github.io/agent-studio/
CRAB [454] Linux, Android 2024 Yes Cross-environment benchmark evaluating agents across mobile and desktop devices, using a graph-based evaluation method to handle multiple correct paths and task flexibility. 120 tasks Step Success Rate, Efficiency Score Action Match https://github.com/crab-benchmark
ScreenSpot [25] iOS, Android, macOS, Windows, Web 2024 No Vision-based GUI benchmark with pre-trained GUI grounding, assessing agents’ ability to interact with GUI elements across mobile, desktop, and web platforms using only screenshots. 1,200 instructions Step Success Rate Action Match https://github.com/njucckevin/SeeClick
@@ 9.7 Cross-Platform Agent Benchmarks
To develop GUI agents capable of operating across multiple platforms, cross-platform benchmarks are essential. These benchmarks challenge agents to adapt to different environments and interfaces, evaluating their versatility and robustness. We provide an overview of benchmarks for cross-platform GUI agents in Tables XXXIX.
@@
Complementing this, CRAB [454] introduces an in vative benchmark by evaluating multimodal language model agents in cross-environment interactions. It uniquely supports seamless multi-device task execution, evaluating agents in scenarios where tasks span both Ubuntu Linux and Android environments. By introducing a graph-based evaluation method that breaks down tasks into sub-goals and accommodates multiple correct paths to completion, CRAB provides a nuanced assessment of planning, decision-making, and adaptability. Metrics such as Completion Ratio, Execution Efficiency, Cost Efficiency, and Success Rate offer comprehensive insights into agent performance.
@@ 9.8 Takeaways
    Towards More Interactive and Realistic Environments: Recent advancements in GUI agent benchmarking emphasize the transition from synthetic scenarios to more interactive and realistic environments. This shift is evident in the use of simulators, Docker containers, and real-world applications to create "live" environments that better mimic genuine user interactions. Such environments t only provide a more accurate assessment of agent capabilities but also pose new challenges in terms of performance and robustness.
@@
    Emphasis on Safety, Privacy, and Compliance: There is a table trend towards evaluating GUI agents on safety, privacy, and compliance metrics. These considerations are increasingly important as agents are integrated into sensitive and regulated domains. Encouraging this trend will help ensure that agents t only perform tasks effectively but also adhere to necessary legal and ethical standards. Future benchmarks should continue to expand on these dimensions, incorporating evaluations that reflect real-world compliance and data security requirements.
The landscape of GUI agent benchmarking is rapidly evolving to meet the demands of increasingly complex and interactive environments. By embracing cross-platform evaluations, fostering human-like interactions, addressing scalability challenges, and prioritizing safety and compliance, the community can pave the way for the next generation of sophisticated GUI agents. Continued in vation and collaboration will be essential in refining benchmarks to ensure they accurately capture the multifaceted capabilities of modern agents, ultimately leading to more intuitive and effective human-computer interactions.
@@ 10.1 GUI Testing
TABLE XL: Overview of GUI-testing with LLM-powered GUI agents (Part I).
@@
TABLE XLI: Overview of GUI-testing with LLM-powered GUI agents (Part II).
@@
UXAgent [467] Usability testing for web design Web Self-designed Simplified HTML representations Standard UI operations Automated usability testing of web applications Enables LLM-powered automated usability testing by simulating thousands of user interactions, collecting both qualitative and quantitative data, and providing researchers with early feedback before real-user studies. https://uxagent.hailab.io
Guardian [468] GUI Testing Mobile Android GPT-3.5 GUI structure, Properties Standard UI operations Autonomously explores mobile applications, interacting with the UI to validate core functionalities. Improves LLM-driven UI testing by offloading planning tasks to an external runtime system. /
Test-Agent [469] GUI Testing Android, iOS, Harmony OS Not Mentioned GUI screenshots, UI structure information Standard UI operations Cross-platform mobile testing Eliminates the need for pre-written test scripts by leveraging LLMs and multimodal perception to generate and execute test cases automatically. /
VLM-Fuzz [470] GUI Testing Android (Mobile) GPT-4o GUI screenshots and UI structure information Standard UI operations, system-level actions Automated Android app testing for detection of crashes and bugs Integrates vision-language reasoning with heuristic-based depth-first search (DFS) to systematically explore complex Android UIs, achieving significantly higher code coverage /
TABLE XLII: Overview of GUI-testing with LLM-powered GUI agents (Part III).
BugCraft [471] Bug Reproduction Windows Computer BugCraft based on GPT-4o GUI screenshots Standard UI operations Automatically reproduces crash bugs in Minecraft by reading user-submitted bug reports, generating structured steps, and executing them to cause a crash First end-to-end framework that automates crash bug reproduction in a complex open-world game (Minecraft) using LLM-driven agents, vision-based UI parsing, and structured action execution https://bugcraft2025.github.io/
ReuseDroid [472] GUI Testing Mobile Android ReuseDroid based on GPT-4o GUI screenshots and widget properties Standard UI operations Migrates GUI test cases between Android apps that share similar functionality but differ in operational logic Leverages visual contexts and dynamic feedback mechanisms to significantly boost migration success rates compared to prior mapping- and LLM-based methods /
SeeAct-ATA and PinATA [473] GUI Testing Web SeeAct [17] GUI structure and DOM Standard UI operations Automates manual end-to-end (E2E) web application testing First open-source attempt to adapt LLM-powered Autonomous Web Agents into Autonomous Test Agents (ATA) for web testing /
GERALLT [474] GUI Testing Desktop (Windows/Linux) GPT-4o GUI screenshots and UI structure information Standard UI operations Finds unintuitive behavior, inconsistencies, and functional errors in GUIs without pre-defined test scripts Pioneers LLM-driven testing on real-world desktop GUI applications (not web or mobile), combining structured GUI parsing with LLM-based control and evaluation https://github.com/DLR-SC/GERALLT
@@
Figure 29 and illustrates the use of an LLM-powered GUI agent to test font size adjustment on Windows OS. With only a natural language test case description, the agent autonomously performs the testing by executing UI operations, navigating through the settings menu, and leveraging its screen understanding capabilities to verify the final outcome of font size adjustment. This approach dramatically reduces the effort required for human or script-based testing. Next, we detail the GUI testing works powered by GUI agents, and first provide an overview Tables XL, XLI and XLII.
@@ 10.1.3 Bug Replay
BugCraft [471] leverages LLM-powered GUI agents to automate bug reproduction in games, specifically targeting the open-ended and complex environment of Minecraft. It employs GPT-4o as the inference engine, integrating textual bug reports, visual GUI understanding through OmniParser [183], and external knowledge from the Minecraft Wiki to generate and execute structured reproduction steps. Actions are carried out via a custom Macro API, enabling robust interaction with both the game’s GUI and environment. BugCraft’s ability to translate unstructured bug descriptions into executable in-game behaviors highlights the strong potential of vision-enhanced LLM agents for advancing software testing and debugging.
@@ 10.2 Virtual Assistants
TABLE XLIII: Overview of virtual assistants with LLM-powered GUI agents (Part I).
@@
WebNav [488] Research Web Gemini 2.0 Flash Thinking Standard UI operations GUI screenshots and DOM Assistive technology for visually impaired users, enabling voice-based navigation of complex websites Combines a ReAct-style reasoning loop, real-time DOM labeling, and voice-driven interaction to support intelligent web navigation for visually impaired users /
TABLE XLIV: Overview of virtual assistants with LLM-powered GUI agents (Part II).
@@
To explore more real-world applications of virtual assistants powered by GUI agents, we provide an overview of advancements across research, open-source initiatives, and production-level applications, as summarized in Table XLIII and XLIV.
@@ 11.1 Privacy Concerns
Privacy is a critical concern uniquely intensified in the context of LLM-powered GUI agents. These agents often require access to sensitive user data—such as screenshots, interaction histories, personal credentials, and confidential documents—to effectively perceive and interact with the GUI environment. In many cases, this data must be transmitted to remote servers for model inference, especially when relying on cloud-based LLMs [537, 538, 539]. Such deployments raise significant privacy risks, including data breaches, unauthorized access, and misuse of personal information. These concerns are further amplified when sensitive inputs are routed through third-party APIs or processed off-device, creating compliance and security vulnerabilities that can deter real-world adoption.
For instance, a GUI agent tasked with managing a user’s email inbox may need to read, classify, and respond to messages containing highly personal or confidential content. Offloading this processing to the cloud introduces risks of exposure, prompting hesitation among users and organizations due to potential privacy violations [426, 540, 541]. Compared to traditional LLM applications, GUI agents operate at a finer granularity of user activity and often require broader system access, making privacy-preserving deployment strategies a critical and domain-specific challenge.
Potential Solutions: To mitigate privacy concerns, future research should focus on enabling on-device inference, where the language model operates directly on the user’s device without uploading personal data [542, 543]. Achieving this requires advancements in model compression techniques [544], on-device optimization [545], and efficient inference algorithms [546] to accommodate the computational limitations of user devices. In addition, frameworks must incorporate data redaction, secure communication channels, and explicit scoping of data usage within the agent’s context. Furthermore, integration with system-level privacy controls and user consent mechanisms (e.g., runtime permission dialogs or sandboxed execution) is essential for deployment in regulated domains.
From the technical perspective, implementing privacy-preserving techniques like federated learning [547], differential privacy [548], and homomorphic encryption [549] can enhance data security while allowing the model to learn from user data. Furthermore, developers of GUI agents should collaborate with privacy policymakers to ensure that user data and privacy are appropriately protected [550]. They should make the data handling processes transparent to users, clearly informing them about what data are being transmitted and how they are used, and obtain explicit user consent [551].
@@ 11.2 Latency, Performance, and Resource Constraints
One challenge that is particularly salient for GUI agents—distinct from general LLM applications is the issue of latency in interactive, multi-step execution environments. Since GUI agents rely on large language models to plan and issue actions, their computational demands can lead to high latency and slow response times, which directly impact user experience [552]. This is especially critical in time-sensitive or interactive scenarios, where delays in action execution can cause user frustration or even trigger unintended system behavior. Unlike single-shot LLM tasks, GUI agents typically operate over extended sequences of steps, making latency cumulative and more disruptive over time. The problem is further amplified in on-device deployments, where computational resources are limited. For example, running an LLM-powered agent within a mobile app may result in sluggish performance or rapid battery depletion, significantly undermining usability on resource-constrained platforms [553, 554, 555]. These concerns are uniquely pronounced in GUI agents due to their need for real-time perception, decision-making, and UI control in dynamic environments [554].
@@ 11.3 Safety and Reliability
The real-world actuation capabilities of GUI agents introduce unique and significant safety and reliability risks beyond those faced by general-purpose LLMs. Because GUI agents can directly manipulate user interfaces—clicking buttons, deleting files, submitting forms, or initiating system-level operations—errors in action generation can have irreversible consequences [562, 539]. These may include data corruption, accidental message dispatches, application crashes, or unauthorized access to sensitive system components [563, 564]. Such risks are compounded by the inherent uncertainty and non-determinism in LLM outputs: agents may hallucinate actions, misinterpret UI contexts, or behave inconsistently across sessions [565, 566, 567, 568, 569]. For example, an agent automating financial transactions could mistakenly execute the wrong transfer, leading to material losses. Furthermore, GUI agents expose a broader attack surface than traditional LLM applications—they are susceptible to black-box adversarial attacks that could manipulate their inputs or exploit their decision policies [570].
Unlike passive language models, GUI agents operate within dynamic software ecosystems where incorrect actions can propagate across applications or escalate into system-wide disruptions. Integration challenges also arise, including compatibility with evolving UI frameworks, user permission boundaries, and software-specific safety constraints, and malicious attacks [571, 572]. These concerns, coupled with the lack of interpretability and formal guarantees, contribute to skepticism and reluctance from users and developers alike. Addressing safety and reliability in GUI agents thus requires not only robust model behavior but also runtime safeguards [573], rollback mechanisms, and interface-aware verification techniques tailored specifically to this interaction paradigm.
@@ 11.4 Human-Agent Interaction
Human-agent interaction introduces distinct challenges in the context of GUI agents, where the agent and user operate within the same dynamic interface. Any user intervention—such as moving the mouse, altering window states, or modifying inputs—can inadvertently interfere with the agent’s ongoing execution, potentially causing conflicts, unintended actions, or breakdowns in task flow [585, 586]. Designing robust collaboration protocols that govern when the agent should yield control, pause execution, or defer to the user is a non-trivial problem specific to GUI-based automation.
Further complicating this interaction is the ambiguity of user instructions. Natural language commands may be vague, under-specified, or context-dependent, leading to misinterpretations or incomplete task plans. GUI agents may also encounter runtime uncertainties—such as unexpected popups, missing inputs, or conflicting UI states—that require them to seek user clarification or feedback [19, 587]. Determining when and how an agent should request user input—whether for disambiguation, permission, or verification—is critical for ensuring both reliability and user trust [588, 67].
This challenge is exemplified in the fabricated scenario shown in Figure 31, where a GUI agent is instructed to send an email to “Tom.” The agent must first prompt the user to log in securely, protecting credentials by avoiding automated input. It then encounters ambiguity when multiple contacts named “Tom” are found, and resolves it by prompting the user to select the intended recipient. Finally, before dispatching the email, the agent requests explicit confirmation, recognizing that email-sending is a non-reversible action with privacy implications [19]. Although the task appears simple, it reflects the complexity of real-world human-GUI agent collaboration, involving privacy preservation, ambiguity resolution, and intentionality confirmation [589]. These are not generic LLM issues, but domain-specific challenges rooted in shared interaction with software interfaces—underscoring the need for new design paradigms around shared control, interruption handling, and proactive clarification in GUI agent systems.
@@
Potential Solutions: Emphasizing user-centered design [590] principles can address user needs and concerns, providing options for customization and control over the agent’s behavior [587]. Equipping agents with the ability to engage in clarification dialogues when user instructions are unclear can enhance task accuracy [591]. Natural language understanding components can detect ambiguity and prompt users for additional information. For instance, the agent could ask, “There are two contacts named John. Do you mean John Smith or John Doe?” Incorporating human-in-the-loop systems allows for human intervention during task execution, enabling users to guide or correct the agent’s decisions when necessary [592]. Developing adaptive interaction models that facilitate seamless collaboration between humans and agents is essential. Additionally, providing transparency and explainability in the agent’s reasoning processes can build user trust and improve cooperation [593, 594, 67].
Figure 32: The Picture-in-Picture interface in UFO2: a virtual desktop window enabling non-disruptive automation. Figure adapted from [330].
Lastly, developing a virtual desktop environment for the agent to operate in—one that connects to the user’s main desktop session without disrupting their workflow, can significantly enhance the user experience (UX) in human-agent interaction. The picture-in-picture mode implemented in UFO2 [330] demonstrates this concept in practice, as illustrated in Figure 32. By allowing the agent to run within a resizable and movable virtualized desktop, users can easily minimize or reposition the agent window as needed. This flexibility improves both the usability and the overall UX of interacting with GUI-based agents.